In electronics, according to Ohm's law, E = IR, where "E" is
voltage (measured in volts), "I" is current (measured in amps), and
"R" is resistance (measured in ohms).
What is the resistance in a circuit having a voltage of 0.02 volts
and a current of 1 milliamp (or 0.001 amps)?

Answer by Alan3354(69443)

R = E/I = 0.02/0.001 ohms